OpenStudy (anonymous):

Every freshman student at a particular college is recquired to take an English aptitude test. A student who passes the examination enrolls in English composition, and a student who fails must enroll in English Fundamentals. In a freshman class of 1240 students, there are more students enrolled in English Fundamentals than in English Composition. However, if 30 more students had passed the test, each course have the same enrollment. How many students are enrolled in each course?

OpenStudy (anonymous):

605 are in Composition 635 are in Fundamentals First subtract 30 to 1240 which results to 1210 then divide in 2 then you get 605 that is for composition.....add 30 to get fundamentals which is 635 :)

OpenStudy (anonymous):

Wait a minute! If 30 passed the test then Composition would become 635 right? And Fundamentals would turn into 635 - 30 = 605. The two courses dont have the same enrollment then. Maybe subtracting 60 from 1240 would resolve this.

OpenStudy (anonymous):

yeah.... i think 650 are in fundamentals and 590 are in composition

OpenStudy (anonymous):

1240 - 60 = 1180 1180/2 = 590 (Compositions) 590 + 60 = 650 (Fundamentals) IF 30 more passed the test then 590 + 30 = 620 (Composition) and 650 - 30 = 620 (Fundamentals) Satisfies the prompt!

OpenStudy (anonymous):

Why subtracting it from 60?

OpenStudy (anonymous):

Never mind I get it now xD